a developer wants to enclose a rectangular lot that borders a city street. if the developer has 248ft of fence and DOES NOT fence the side bordering the street, what is the largest area that can be enclosed?
Answer by checkley77(12844)

TO SOLVE THESE PROBLEMS JUST DIVIDE THE TOTAL LENGTH BY 3.
248/3=62 THIS IS THE SHORT SIDES OF THE FENCE.
THE REMAINING FENCE (248-2*62)=248-124=124 IS THE LONG SINGLE SIDE.
PROOF;
62*124=7,688 SQUARE FT. IS THE LARGEST AREA.
TRY A SQUARE BY DIVIDING THE FENCE LENGTH BY 3.
248/3=82.67
NOW SQUARE 62.67^2=6,834 SQUARE FT.